40, The polynomial x6 – 8 is divisible by?

A.     X-√2

B.     X2 + 2

C.     X3 -3

D.     None of the above


1
Expert's answer
2021-09-02T00:05:53-0400

"(x^6-8)÷(x-\sqrt{2})"


"\frac{x^6-8}{x-\sqrt{2}}"


using "a^2-b^3=(a-b)(a+b)," factor the expression


"\frac{(x^3-\sqrt8)×(x^3+\sqrt{8}}{x-\sqrt{2}}"


Simplify the expressions

"\frac{(x^3-2\sqrt2)×(x^3+2\sqrt{2}}{x-\sqrt{2}}"


Factor the expression

"\frac{(x-\sqrt2)×(x^2+\sqrt{2}+2)×(x^3+2\sqrt{2})}{x-\sqrt{2}}"


Reduce the fraction

"(x^2+\sqrt{2}x+2)×(x^3+2\sqrt{2})"


"x^5+2\sqrt{2}x^2+\sqrt{2}x^4+4x+2x^3+4\sqrt{2}"



Hence the polynomial "x-\sqrt2"